# Cross-positive linear maps, positive polynomials and sums of squares

### v1

@@ 3. Nichtnegativstellensätze for the case $n=3$
In this section we will establish, in the case $n=3$, some certificates of global nonnegativity for the form $q_{A}({\tt x},\alpha)$ of (2.4) or nonnegativity for $p_{A}({\tt x},{\tt y})$ of (1.3) on $V(I)$. By Proposition 2.4, this yields certificates for a linear map $A:M_{3}(\mathbb{R})_{\textrm{sym}}\to M_{3}(\mathbb{R})_{\textrm{sym}}$ to be cross-positive.
@@
and since $p_{A}(({\tt x}_{1},{\tt x}_{2}),({\tt x}_{2},-{\tt x}_{1}))$ is a quartic form, it folllows by [Hil88] that $q_{A}$ is nonnegative if and only if it is a sum of squares.
For a matrix polynomial $A({\tt x})\in M_{m}(\mathbb{R}[{\tt x}])$ we denote by $\operatorname{tr}(A({\tt x}))$ its trace, i.e., the sum of the diagonal entries. For a ring $R$ we denote by $\sum M_{m}(R)^{2}$ the set of all finite sums of the expressions of the form $G^{T}G$, where $G\in M_{m}(R)$. Every element of $\sum M_{m}(R)^{2}$ is called a sum of squares (sos) matrix polynomial. We say a symmetric matrix polynomial $A({\tt x})\in M_{m}(\mathbb{R}[{\tt x}])_{\textrm{sym}}$ is positive semidefinite (psd) in $x\in\mathbb{R}^{n}$ if $v^{T}A(x)v\geq 0$ for every $v\in\mathbb{R}^{m}$, and write $A(x)\succeq 0$. We call $A({\tt x})\in M_{m}(\mathbb{R}[{\tt x}])_{\textrm{sym}}$ psd if it is psd in every $x\in\mathbb{R}^{n}$.
In this paragraph we connect, for any $n\in\mathbb{N}$, a global nonnegativity of $q_{A}({\tt x},\alpha)$ with a positive semidefiniteness of a certain matrix polynomial. We denote by $\mathbb{R}[{\tt x}]_{\hom}$ the set of homogeneous real poynomials in ${\tt x}$. Since $q_{A}({\tt x},\alpha)$ is a quadratic form in $\alpha$ with coefficients from $\mathbb{R}[{\tt x}]_{\hom}$, we can associate to it a symmetric matrix polynomial $Q_{A}\in M_{n-1}(\mathbb{R}[{\tt x}]_{\hom})$ such that
@@
The polynomial $q_{A}({\tt x},\alpha)$ is globally nonnegative iff $Q_{A}({\tt x})$ is positive semidefinite for all ${\tt x}$.
@@
[…]
@@
The nontrivial implication is ()($\Rightarrow$)( ⇒ ). We write $Q=\left[\begin{array}[]{cc}a&b\\ b&c\end{array}\right]$. It is easy to check that
@@
where $V=\frac{1}{2}\left[\begin{array}[]{cc}1&i\\ i&1\end{array}\right]$. By (3.8), we have
(3.16) $\operatorname{tr}(Q)^{4}\;Q=\widetilde{V}\left[\begin{array}[]{cc}\operatorname{tr}(Q)^{3}&0\\ 0&\operatorname{tr}(Q)(\operatorname{tr}(Q)^{2}-dd^{\ast})\end{array}\right]\widetilde{V}^{\ast}$
where $\widetilde{V}=V\left[\begin{array}[]{cc}\operatorname{tr}(Q)&0\\ d^{\ast}&\operatorname{tr}(Q)\end{array}\right]$ and $d:=i(a-c)+2b$. A short computation shows that $\operatorname{tr}(Q)^{2}-dd^{\ast}=4\det Q$. Since the left hand side of (3.16) belongs to $M_{2}(\mathbb{R}[x]_{\hom})$, the right hand side equals
@@
where $V_{1},V_{2}\in M_{2}(\mathbb{R}[x])$ are the real and imaginary parts of $\widetilde{V}$. Now iv follows by [Sce06, Corollary 3.12] since there exists $N\in\mathbb{N}$ large enough such that both forms on the diagonal of $D$ multiplied by $({\tt x}_{1}^{2}+{\tt x}_{2}^{2}+{\tt x}_{3}^{2})^{N}$ are sum of squares of forms. ∎
@@
If $Q$ in Theorem 3.3 is quartic (for example, $Q=Q_{A}$), then $\operatorname{tr}Q$ is a ternary quartic. Thus it is a sum of three squares by [Hil88]. So in that case the exponent $N$ in iv of Theorem 3.3 depends only on $\det Q$ which is of degree 8. By [Hil93] there is a positive form $q$ of degree $4$ such that $q\det Q$ is a sum of squares of three forms. Moreover, $q^{2}\det Q$ is a sum of squares of four forms [Lan86]. See also [Rez05, p. 2830].
@@
Claim. There exists an orthogonal matrix $U\in M_{2}(\mathbb{R})$ such that, denoting $UQU^{T}=\left[\begin{array}[]{cc}\widetilde{a}({\tt x})&\widetilde{b}({\tt x})\\ \widetilde{b}({\tt x})&\widetilde{c}({\tt x})\end{array}\right]$, we have that $\widetilde{a}(x_{0})\neq 0$.
@@
It follows that $d_{1}\geq 0$, $d_{2}\geq 0$ on $\mathbb{R}^{3}$. By [Sce06, Theorem 3.2], $d_{1}$ and $d_{2}$ belong to $\sum\mathbb{R}[{\tt x}]^{2}+J$. This concludes the proof. ∎
@@
Let $R:=\mathbb{R}[{\tt x}]/J$ be a quotient ring and let $\Phi:R\rightarrow C(V(J),\mathbb{R})$ be the natural map, i.e., $\Phi(\check{f})=f|_{V(J)},$ where $f\in\mathbb{R}[{\tt x}]$, $\check{f}=f+J$ and the variety $V(J)$ is the set $\left\{x\in\mathbb{R}^{3}\colon\sum_{i=1}^{3}x_{i}^{2}=1\right\}.$ Let
@@
be an ideal in $\mathbb{R}[{\tt x}]/J$. If $L$ was a proper ideal, then all its elements would have a common zero $x_{0}\in\{x\in\mathbb{C}^{3}\colon\sum_{i=1}^{3}x_{i}^{2}=1\}$. By Proposition 3.6, there exists $h\in\mathbb{R}[{\tt x}]$ such that $h(x_{0})\neq 0$ and $\check{h}^{2}\in L$. Hence $L$ is not a proper ideal and thus $L=\mathbb{R}[{\tt x}]/J$. In particular, there are $\check{h}_{1}^{2},\ldots,\check{h}_{k}^{2}\in L$ such that $1+J\in\langle\check{h}_{1}^{2},\ldots,\check{h}_{k}^{2}\rangle$. By [Sce06, Proposition 2.7], there exist $s_{1},\ldots,s_{k}\in\mathbb{R}[{\tt x}]$ with $s_{j}>0$ on $V(J)$ such that $\sum_{j=1}^{k}s_{j}h_{j}^{2}\in 1+J$. By [Sce06, Theorem 3.2], $s_{j}\in\sum\mathbb{R}[{\tt x}]^{2}+J$. Hence
@@
The following lemma, which holds in any dimension $n$, gives a special sufficient condition for a biquadratic form $p_{A}$ that is nonnegative on $V(I)$ to be a sum of a globally nonnegative biquadratic form and an element of $I$. Using the language of [KOŠT15], a cross-positive map $A$ satisfying this condition is not exotic. The lemma will be then used in Corollary 3.10 to establish the second main result of this section: a certificate for nonnegativity of $p_{A}$ on $V(I)$ in the case $n=3$ when $Q_{A}$ has a nontrivial real zero.
@@
For each $i=1,\ldots,n$ let $e_{i}$ be the $i$-th element of the standard basis of $\mathbb{R}^{n}$, i.e., the vector with 1 on the $i$-th component and 0 elsewhere. We denote by $E_{ij}:=e_{i}e_{j}^{T}$ the standard $n\times n$ matrix units.
@@
In the proof of Lemma 3.7 and Corollary 3.10 below we will use the following action of $\mathrm{GL}_{n}$ on the set of cross-positive linear maps $A:M_{n}(\mathbb{R})_{\mathrm{sym}}\to M_{n}(\mathbb{R})_{\mathrm{sym}}$:
@@
    There exists a matrix $C\in M_{n}(\mathbb{R})$ such that the map $B:M_{n}(\mathbb{R})_{\textrm{sym}}\to M_{n}(\mathbb{R})_{\textrm{sym}}$, defined by $B(X):=A(X)-CX-XC^{T}$, satisfies the following conditions:
@@
It remains to check (3). We write $x=\lambda e_{1}+v$ and $y=\mu e_{1}+w$ for some $v,w\perp e_{1}$ and some $\lambda,\mu\in\mathbb{R}$. Using (3.20)–(3.22) we have that $B(e_{1}e_{i}^{T}+e_{i}e_{1}^{T})=0$ for each $i=1,\ldots,n$ and thus $B(xx^{T})=B(vv^{T})=B((\nu e_{1}+v)(\nu e_{1}+v)^{T})$ for each $\nu\in\mathbb{R}$. If $\mu\neq 0$, then $y^{T}(v-\frac{w^{T}v}{\mu}e_{1})=0$, therefore cross-positivity of $B$ implies
@@
As a consequence of Lemma 3.7 it follows that testing cross-positivity of a linear map is a NP-hard problem, since this is true for testing positivity of a biquadratic form, see, e.g., [LNQY10, Theorem 2.2]. Indeed, let $A\colon M_{n}(\mathbb{R})_{\textrm{sym}}\to M_{n}(\mathbb{R})_{\textrm{sym}}$ be an arbitrary linear map. Let $D\colon M_{n+1}(\mathbb{R})_{\textrm{sym}}\to M_{n}(\mathbb{R})_{\textrm{sym}}$ be the map that deletes the first row and column and let $B\colon M_{n+1}(\mathbb{R})_{\textrm{sym}}\to M_{n+1}(\mathbb{R})_{\textrm{sym}}$ be defined by $B(X)=\left[\begin{array}[]{cc}0&0\\ 0&A(D(X))\end{array}\right]$. It is clear that the map $A$ is positive iff $B$ is. However, $B(E_{11})=0$ and $B(E_{1i}+E_{i1})=0$ for $i=2,\ldots,n+1$, hence the proof of Lemma 3.7 implies that $B$ is positive iff it is cross-positive. Thus the problem of testing whether $B$ is cross-positive is from a computational complexity viewpoint at least as hard as checking whether $A$ is positive, implying that testing cross-positivity of a map is NP-hard.
@@
By symmetry we can assume that there exist $v_{0}\in\mathbb{R}^{3}\backslash\{0\}$ and linearly independent vectors $w_{1},w_{2}\perp v_{0}$ such that $p(v_{0},w_{1})=p(v_{0},w_{2})=0$. As $p$ is nonnegative on $V(I)$ it is equal to $p_{A}$ for some cross-positive map $A\colon M_{3}(\mathbb{R})_{\textrm{sym}}\to M_{3}(\mathbb{R})_{\textrm{sym}}$. By Remark 3.8 we may assume that $v_{0}=e_{1}$. By the assumption of the corollary the quadratic form $(\lambda,\mu)\mapsto p_{A}(e_{1},\lambda w_{1}+\mu w_{2})$ is positive semidefinite with zero coefficients at $\lambda^{2}$ and at $\mu^{2}$. Consequently, $p_{A}(e_{1},\lambda w_{1}+\mu w_{2})=0$ for all $\lambda,\mu\in\mathbb{R}$, or equivalently, $w^{T}A(e_{1}e_{1}^{T})w=0$ for each $w\perp e_{1}$. By Lemma 3.7 there exists $C\in M_{3}(\mathbb{R})$ such that the map $B\colon M_{3}(\mathbb{R})_{\textrm{sym}}\to M_{3}(\mathbb{R})_{\textrm{sym}}$, defined by $B(X)=A(X)-CX-XC^{T}$, is positive. The biquadratic form $p_{B}$ satisfies the assumptions of [Qua15, Lemma 4.2], so it is a sum of squares of bilinear forms. As $p_{A}-p_{B}\in I$ by the construction of $B$, this proves the corollary.∎
4. Blekhermann type volume estimates
In this section we quantify the gap between cross-positive and completely cross-positive maps by extending the estimates on the volumes of compact sections of the cones of nonnegative biforms established in [KMŠZ19] to nonnegative biforms on the variety $V(I)$. The proofs are analoguous to those in [KMŠZ19] and are inspired by [Ble06, BB05].

### v2

@@ 3. Nichtnegativstellensätze for the case $n=3$
In this section we will establish, in the case $n=3$, some certificates of global nonnegativity for the form $q_{A}({\tt x},\alpha)$ of (2.4) or nonnegativity for $p_{A}({\tt x},{\tt y})$ of (1.3) on $V(I)$. By Proposition 2.5, this yields certificates for a $\ast$-linear map $A:M_{3}(\mathbb{R})\to M_{3}(\mathbb{R})$ to be cross-positive.
@@
and since $p_{A}(({\tt x}_{1},{\tt x}_{2}),({\tt x}_{2},-{\tt x}_{1}))$ is a quartic form, it follows by [19] that $q_{A}$ is nonnegative if and only if it is a sum of squares.
For a matrix polynomial $A({\tt x})\in M_{m}(\mathbb{R}[{\tt x}])$ we denote by $\operatorname{tr}(A({\tt x}))$ its trace, i.e., the sum of the diagonal entries. For a ring $R$ we denote by $\sum M_{m}(R)^{2}$ the set of all finite sums of the expressions of the form $G^{T}G$, where $G\in M_{m}(R)$. Every element of $\sum M_{m}(R)^{2}$ is a sum of squares (sos) matrix polynomial. We say a symmetric matrix polynomial $A({\tt x})\in M_{m}(\mathbb{R}[{\tt x}])_{\textrm{sym}}$ is positive semidefinite (psd) in $x\in\mathbb{R}^{n}$ if $v^{T}A(x)v\geq 0$ for every $v\in\mathbb{R}^{m}$, and write $A(x)\succeq 0$. We call $A({\tt x})\in M_{m}(\mathbb{R}[{\tt x}])_{\textrm{sym}}$ psd if it is psd in every $x\in\mathbb{R}^{n}$.
In this paragraph we connect, for every $n\in\mathbb{N}$, global nonnegativity of $q_{A}({\tt x},\alpha)$ with positive semidefiniteness of a certain matrix polynomial. We denote by $\mathbb{R}[{\tt x}]_{\hom}$ the set of homogeneous real polynomials in $\tt x$}. Since $q_{A}({\tt x},\alpha)$ is a quadratic form in $\alpha$ with coefficients from $\mathbb{R}[{\tt x}]_{\hom}$, we can associate to it a symmetric matrix polynomial $Q_{A}\in M_{n-1}(\mathbb{R}[{\tt x}]_{\hom})$ such that
@@
The polynomial $q_{A}({\tt x},\alpha)$ is globally nonnegative if and only if $Q_{A}({\tt x})$ is positive semidefinite for all $\tt x$}.
@@
Remark 3.3.
Note that in the case $n=3$, Proposition 3.2 implies that the parameterization (2.3) leads to the reduction of the problem of certifying cross-positivity of the map $A$ to certifying positivity of the $2\times 2$ matrix polynomial $Q_{A}$. Under the assumption that $Q_{A}$ does not vanish in any point of $\mathbb{R}^{3}$ we establish such a certificate in Theorem 3.4 below.
@@
For $Q=\left[\begin{array}[]{cc}a&b\\ b^{\ast}&c\end{array}\right]\in M_{2}(\mathbb{C}[{\tt x}])_{\textrm{her}}$ the following equalities hold:
@@
The nontrivial implication is $(\Rightarrow)$. We write $Q=\left[\begin{array}[]{cc}a&b\\ b&c\end{array}\right]$. It is easy to check that
@@
where $V=\frac{1}{2}\left[\begin{array}[]{cc}1&i\\ i&1\end{array}\right]$. By (3.8),
(3.16) $\operatorname{tr}(Q)^{4}\;Q=\widetilde{V}\left[\begin{array}[]{cc}\operatorname{tr}(Q)^{3}&0\\ 0&\operatorname{tr}(Q)(\operatorname{tr}(Q)^{2}-dd^{\ast})\end{array}\right]\widetilde{V}^{\ast}$,
where $\widetilde{V}=V\left[\begin{array}[]{cc}\operatorname{tr}(Q)&0\\ d^{\ast}&\operatorname{tr}(Q)\end{array}\right]$ and d:=i(ac)+2bd:=i(a-c)+2b. A computation shows $\operatorname{tr}(Q)^{2}-dd^{\ast}=4\det Q$. Since the left hand side of (3.16) belongs to $M_{2}(\mathbb{R}[x]_{\hom})$, the right hand side equals
@@
where $V_{1},V_{2}\in M_{2}(\mathbb{R}[x])$ are the real and imaginary parts of $\widetilde{V}$. Now iv follows by [36, Corollary 3.12] since there exists $N\in\mathbb{N}$ large enough such that each form on the diagonal of $D$ multiplied by $({\tt x}_{1}^{2}+{\tt x}_{2}^{2}+{\tt x}_{3}^{2})^{N}$ is a sum of squares of forms. ∎
@@
If $Q$ in Theorem 3.4 is quartic (for example, $Q=Q_{A}$), then $\operatorname{tr}Q$ is a ternary quartic. Thus it is a sum of three squares by [19]. So in that case the exponent $N$ in iv of Theorem 3.4 depends only on $\det Q$ which is of degree 8. By [20] there is a positive form $q$ of degree $4$ such that $q\det Q$ is a sum of squares of three forms. Moreover, $q^{2}\det Q$ is a sum of squares of four forms [24]. See also [35, p. 2830].
@@
Claim. There exists an orthogonal matrix $U\in M_{2}(\mathbb{R})$ such that, denoting $UQU^{T}=\left[\begin{array}[]{cc}\widetilde{a}({\tt x})&\widetilde{b}({\tt x})\\ \widetilde{b}({\tt x})&\widetilde{c}({\tt x})\end{array}\right]$, $\widetilde{a}(x_{0})\neq 0$.
@@
It follows that $d_{1}\geq 0$, $d_{2}\geq 0$ on $\mathbb{R}^{3}$. By [36, Theorem 3.2], $d_{1}$ and $d_{2}$ belong to $\sum\mathbb{R}[{\tt x}]^{2}+J$. This concludes the proof. ∎
@@
[…]
@@
be an ideal in $\mathbb{R}[{\tt x}]/J$. If $L$ were a proper ideal, then all its elements would have a common zero $x_{0}\in\{x\in\mathbb{C}^{3}\colon\sum_{i=1}^{3}x_{i}^{2}=1\}$. By Proposition 3.7, there exists $h\in\mathbb{R}[{\tt x}]$ such that $h(x_{0})\neq 0$ and $\check{h}^{2}\in L$. Hence $L$ is not a proper ideal and thus $L=\mathbb{R}[{\tt x}]/J$. In particular, there exist $\check{h}_{1}^{2},\ldots,\check{h}_{k}^{2}\in L$ such that $1+J\in\langle\check{h}_{1}^{2},\ldots,\check{h}_{k}^{2}\rangle$. By [36, Proposition 2.7], there exist $s_{1},\ldots,s_{k}\in\mathbb{R}[{\tt x}]$ with $s_{j}>0$ on $V(J)$ such that $\sum_{j=1}^{k}s_{j}h_{j}^{2}\in 1+J$. By [36, Theorem 3.2], $s_{j}\in\sum\mathbb{R}[{\tt x}]^{2}+J$. Hence
@@
The following lemma, which holds in every dimension $n$, gives a special sufficient condition for a biquadratic form $p_{A}$ that is nonnegative on $V(I)$ to be a sum of a globally nonnegative biquadratic form and an element of $I$. Using the language of [23], a cross-positive map $A$ satisfying this condition is not exotic. The lemma will be used in Corollary 3.11 to establish the second main result of this section: a certificate for nonnegativity of $p_{A}$ on $V(I)$ in the case $n=3$ when $Q_{A}$ has a nontrivial real zero.
@@
Let $A\colon M_{n}(\mathbb{R})\to M_{n}(\mathbb{R})$ be a cross-positive map and assume that there exists a nonzero vector $x_{0}\in\mathbb{R}^{n}$ such that
@@
For each $i=1,\ldots,n$ let $e_{i}$ be the $i$-th element of the standard basis of $\mathbb{R}^{n}$, i.e., the vector with 1 in the $i$-th component and 0 elsewhere. We denote by $E_{ij}:=e_{i}e_{j}^{T}$ the standard $n\times n$ matrix units.
@@
In the proof of Lemma 3.8 and Corollary 3.11 below we will use the following action of $\mathrm{GL}_{n}$ on the set of cross-positive linear maps $A:M_{n}(\mathbb{R})\to M_{n}(\mathbb{R})$:
@@
    There exists a matrix $C\in M_{n}(\mathbb{R})$ such that the map $B:M_{n}(\mathbb{R})\to M_{n}(\mathbb{R})$, defined by $B(X):=A(X)-CX-XC^{T}$, satisfies the following conditions:
@@
It remains to check (3). We write $x=\lambda e_{1}+v$ and $y=\mu e_{1}+w$ for some $v,w\perp e_{1}$ and some $\lambda,\mu\in\mathbb{R}$. Using (3.20)–(3.22) we see that $B(e_{1}e_{i}^{T}+e_{i}e_{1}^{T})=0$ for each $i=1,\ldots,n$ and thus $B(xx^{T})=B(vv^{T})=B((\nu e_{1}+v)(\nu e_{1}+v)^{T})$ for each $\nu\in\mathbb{R}$. If $\mu\neq 0$, then $y^{T}(v-\frac{w^{T}v}{\mu}e_{1})=0$, therefore cross-positivity of $B$ implies
@@
As a consequence of Lemma 3.8 it follows that testing cross-positivity of a linear map is a NP-hard problem, since this is true for testing positivity of a biquadratic form, see, e.g., [26, Theorem 2.2]. Indeed, let $A\colon M_{n}(\mathbb{R})\to M_{n}(\mathbb{R})$ be an arbitrary $\ast$-linear map. Let $D\colon M_{n+1}(\mathbb{R})\to M_{n}(\mathbb{R})$ be the map that deletes the first row and column and let $B\colon M_{n+1}(\mathbb{R})\to M_{n+1}(\mathbb{R})$ be defined by $B(X)=\left[\begin{array}[]{cc}0&0\\ 0&A(D(X))\end{array}\right]$. It is clear that the map $A$ is positive if and only if $B$ is. However, $B(E_{11})=0$ and $B(E_{1i}+E_{i1})=0$ for $i=2,\ldots,n+1$, hence the proof of Lemma 3.8 implies that $B$ is positive if and only if it is cross-positive. Thus the problem of testing whether $B$ is cross-positive is from a computational complexity viewpoint at least as hard as checking whether $A$ is positive, implying that testing cross-positivity of a map is NP-hard.
@@
[…]
4. Blekherman type volume estimates
In this section we quantify the gap between cross-positive and completely cross-positive maps by extending the estimates on the volumes of compact sections of the cones of nonnegative biforms established in [22] to nonnegative biforms on the variety $V(I)$. The proofs are analogous to those in [22] and are inspired by [5, 3].